For a certain rectangle, the ratio of the shorter side to the longer side is 4 inches to 9 inches. How long is the shorter side, in feet, if the longer side is half a foot long?
Answer by jim_thompson5910(35256)

4/9 = x/0.5

4*0.5 = 9x

2 = 9x

2/9 = x

x = 2/9

x = 0.222222

So the shorter side is 0.2222 ft or about 0.222*12 = 2.664 inches long.